# Approximate Nash Equilibria Algorithms for Network Design Games

### 05 Dec 2023 v1

@@ Abstract
We consider a weighted network design game, where selfish players chooses paths in a network to minimize their cost. The cost function of each edge in the network is affine linear, namely $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, where $a_{e},b_{e}>0$ are only related to the edge $e$ and $W_{e}$ is the total weight of the players that choose a path containing edge $e$. We first show the existence of $\alpha$-approximate Nash equilibrium and prove the upper bound of $\alpha$ is $O(log_{2}(W))$, where $W$ is the sum of the weight of all players. Furthermore, considering that compute the $\alpha$-approximate Nash equilibrium is PLS-complete, we assume that $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables on $[0,1]$. In this case, we show that $\epsilon$-better response dynamics can compute the $\alpha$-approximate Nash Equilibrium in polynomial time by proving the expected number of iterations is polynomial in $\frac{1}{\epsilon},\phi$, the number of players and the number of edges in the network.
@@
Approximate Nash equilibrium Affine linear cost function Better response dynamics
@@ 1 Introduction
Along with the development of many networks including the Internet, the number of agents using networks is increasing, all of whom are motivated by self-interest and optimize an individual objective function. This naturally leads us to study the behavior of these non-cooperative agents in the game. Moreover, the outcome of the interactions of selfish agents is also worthy of analysis, and the stable outcomes are called Nash equilibrium. Furthermore, the study of Nash equilibrium gives rise to a number of new issues. Specifically, since the overall performance of the system brought about by the selfish agents making their own decisions may be worse than that brought about by centralized authority decision-making, we need an indicator to measure the degree to which the global performance of the system has deteriorated. Papadimitriou [1] first proposed using the term price of anarchy to refer to the deterioration of system performance. Moreover, there is more than one Nash equilibrium in some types of games, but the price of anarchy only studies the deterioration in the worst Nash equilibrium. Then Anshelevich et al. [3] introduced the term price of stability to refer to the ratio of the cost of best Nash equilibrium to the optimum network cost.
On the other hand, computing a Nash equilibrium is essential when the game is given from a computational perspective. However, Fabrikant et al. [5] proved that computing a Nash equilibrium of a congestion game is PLS-complete. Furthermore, Ackermann et al. [6] showed that the PLS-completeness holds in network design game and other kinds of games. After that, scholars have turned their direction to computing approximate Nash equilibrium. However, Skopalik et al. [7] proved that computing an $\alpha$-PNE is PLS-complete for any given constant $\alpha$. Therefore, it is necessary to further restrict the game ensure that there is a method to compute the (approximate) Nash equilibrium in polynomial time.
@@ 1.1 Related Works
1.1.1 Network Design Game
Papadimitriou [1] first proposed using algorithmic ideas and game theory to analyze the cost of network. Anshelevich et al. [3] proposed the network design with Shapley cost sharing, which uses a cost-sharing mechanism. However Anshelevich et al. [3] only proved few results when we players are weighted. H. Chen and T. Roughgarden [8] showed that weighted Shapley network design game does not admit a Nash equilibrium. Moreover, they proved the existence of $\alpha$-approximate Nash equilibrium and gives a upper bound of $\alpha$. However, no previous study has been found for the case where the cost function is a linear function to the total weight of the players that choose the edge.
@@ 1.2 Our contributions
Our model extend the cost function in the model proposed by Anshelevich et al. [3], which is a constant $c_{e}$ for each edge $e$ in the network. In our model, the cost function of each edge in the network is affine linear, namely $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, where $a_{e},b_{e}>0$ are only related to the edge $e$ and $W_{e}$ is the total weight of the players that choose a path containing edge $e$. We show the existence of $\alpha$-approximate Nash equilibrium and prove the upper bound of $\alpha$ is $O(log_{2}(W))$, where $W$ is the sum of the weight of all players.
Furthermore, as Skopalik et al. [7] proved that computing an $\alpha$-PNE is PLS-complete for any given constant $\alpha$ in congestion games, our model assume that $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables on $[0,1]$, which means that the cumulative distribution function of $a_{e},b_{e}$ is upper bounded by a constant $\phi\geq 1$. In this case, we show that $\epsilon$-better response dynamics can compute the $\alpha$-approximate Nash Equilibrium in polynomial time by proving the expected number of iterations is polynomial in $\frac{1}{\epsilon},\phi$, the number of players $n$ and the number of edges $m$ in the network.
@@ 2 Model and Preliminaries
We represent an arbitrary weighted network design game $\Gamma$ by tuple $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ with components defined below.
  • P1: $G=(V,E)$ is a directed graph, where $V$ is the vertex set of $G$ and $E$ is the edge set of $G$. Here we further define $m=|E|$, which is the number of edges in graph $G$.
  • P2: $\mathcal{N}$ is a collection of $n$ different players. Moreover, each player $i\in\mathcal{N}$ is identified with a source-sink vertex pair $(s_{i},t_{i})$.
  • P3: $\mathcal{P}=(\mathcal{P}_{i})_{i\in\mathcal{N}}$, where $\mathcal{P}_{i}$ denote the set of simple $s_{i}-t_{i}$ paths. Furthermore, each player $i$ chooses a path $P_{i}\in\mathcal{P}_{i}$ from its source to its sink.
  • P4: $\mathcal{C}=(c_{e})_{e\in E}$, where $c_{e}$ is the non-negative cost of the edge $e$.
@@
All players are selfish, and each of them chooses a path independently. Then it results in a pure strategy profile $P=(P_{i})_{i\in\mathcal{N}}$, where $P_{i}$ denotes the simple $s_{i}-t_{i}$ path used by player $i\in\mathcal{N}$. For a given profile and an edge $e\in E$, we define $U_{e}(P):=\{i|i\in\mathcal{N},e\in P_{i}\}$ as the “user” of edge $e$. Moreover, we define $W_{e}=\sum_{i\in U_{e}(P)}w_{i}$ as the total weight of the players that choose a path containing edge $e$. Furthermore, for a given profile and a player $i$, the cost share $c^{i}_{e}$ of an edge $e\in E$ is $c_{e}\cdot w_{i}/W_{e}$. Then the cost to player $i$ in an profile is the sum of its cost shares: $c_{i}(P)=\sum_{e\in P_{i}}c^{i}_{e}$. In addition, the cost $c(P)$ of a profile $P$ is defined as $\sum_{i\in\mathcal{N}}c_{i}(P)$, which can be also written as $\sum_{e\in\cup_{i\in\mathcal{N}}P_{i}}c_{e}$.
@@
(Pure Strategy Nash Equilibrium) Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game. We call a profile $P=(P_{i})_{i\in\mathcal{N}}$ a pure strategy Nash equilibrium(PNE) if, for each $i$, $P_{i}$ minimizes $c_{i}(P)$ over all paths in $\mathcal{P}_{i}$ while keeping $P_{j}$ fixed for $j\neq i$. That is:
@@
(Exact Potential Function) Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game. We call a real-valued function $\Phi:\prod_{i\in\mathcal{N}}\mathcal{P}_{i}\rightarrow[0,\infty)$ an (exact) potential function if
@@
However, in some cases, the game may not admit a PNE. Then we have to consider a relaxed notion of $\alpha$-approximate pure strategy Nash equilibrium instead.
@@
($\alpha$-Approximate Pure Strategy Nash Equilibrium) Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game and an arbitrary constant $\alpha\geq 1$. We call a profile $P=(P_{i})_{i\in\mathcal{N}}$ an $\alpha$-approximate pure strategy Nash equilibrium ($\alpha$-APNE) if
@@
In fact, Inequality (3) states that a unilateral change of strategy reduces the cost at a rate of at most $\frac{c_{i}(P_{i},P_{-i})-c_{i}(P^{\prime}_{i},P_{-i})}{c_{i}(P_{i},P_{-i})}\leq 1-\frac{1}{\alpha}$ in an $\alpha$-APNE. Moreover, we call a deviation $\alpha$-improving if the deviation decreases the cost incurred by the player by at least an $\alpha$ multiplicative factor. Therefore, $\alpha$-APNE are profiles from which no $\alpha$-improving deviations exist. We further point out that the smaller the constant $\alpha$ is, the closer the profile $P$ would approximate a PNE.
@@
($\alpha$-Approximate Potential Function) Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game. For any player $i\in\mathcal{N}$, $P_{i},P^{\prime}_{i}\in\mathcal{P}_{i}$ satisfies $P_{i}\rightarrow P^{\prime}_{i}$ is an $\alpha$-improving deviation. We call a real-valued function $\Phi:\prod_{i\in\mathcal{N}}\mathcal{P}_{i}\rightarrow[0,\infty)$ an $\alpha$-approximate potential function if
@@
$\phi$-smooth random variable. The probability density function of each random variable is up to $\phi$($\phi\geq 1$).
@@ 3 Existence of $\alpha$-APNE
Anshelevich et al. [3] proved that the unweighted network design game admits a PNE, and pointed out that weighted Shapley network design game does not admit a PNE. Therefore, it is essential to turn the attention to $\alpha$-APNE.
@@
Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game. If the cost function is affine linear, namely $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, where $a_{e},b_{e}>0$ are only related to the edge $e$ and $W_{e}$ is the total weight of the players that choose a path containing edge $e$, then the game admits an $\alpha$-APNE, where $\alpha=O(log_{2}(W))$.
Before proving the theorem above, we first present a lemma.
Lemma 1
Let $x,y\in[1,\infty)$, then we have:

### 17 Dec 2023 v2

# Approximate Nash Equilibria Algorithms for Shapley Network Design Games

@@ Abstract
We consider a weighted Shapley network design game, where selfish players choose paths in a network to minimize their cost. The cost function of each edge in the network is affine linear with respect to the sum of weights of the players who choose the edge. We first show the existence of $\alpha$-approximate pure Nash equilibrium by constructing a potential function and establish an upper bound $O(\log_{2}(W))$ of $\alpha$, where $W$ is the sum of the weight of all players. Furthermore, we assume that the coefficients of the cost function (affine linear function) of the edge all are $\phi$-smooth random variables on $[0,1]$. In this case, we show that $\epsilon$-best response dynamics can compute the $(1+\epsilon)\alpha$-approximate pure Nash equilibrium ($\epsilon$ is a positive constant close to 0) in polynomial time by proving the expected number of iterations is polynomial in $\frac{1}{\epsilon},\phi$, the number of players and the number of edges in the network.
@@
Approximate Nash equilibrium Affine linear cost function Best response dynamics Potential function
@@ 1 Introduction
Along with the development of many networks including the Internet, the number of players using networks is increasing, all driven by self-interest and optimize an individual objective function. This naturally leads us to study the behavior of these non-cooperative players in the game, including [5, 6, 15, 16]. Specifically, the game is defined on a directed network and each player needs to choose a path from a source vertex to a sink vertex. Each edge in the network has a cost and the players using the edge share the cost equally or according to their weights. All players want to reduce personal cost and choose the path simultaneously, forming a profile of the game. The stable profiles are called Nash equilibrium, in which no player has an incentive to unilaterally change strategy. Rosenthal [8] first studied the equilibrium of the game on a network and proved the existence of pure Nash equilibrium (PNE) in the game. Then Anshelevich et al. [3] named the game network design with Shapley cost sharing, also known as Shapley network design game or fair cost sharing game since it applies the cost sharing mechanism.
@@
Furthermore, the study of PNE gives rise to a number of new issues. Specifically, since the overall performance of the system brought about by the selfish players making their own decisions may be worse than that brought about by centralized authority decision-making, we need an indicator to measure the degree to which the global performance of the system has deteriorated. Papadimitriou [2] first proposed using the term price of anarchy to refer to the deterioration of system performance. Moreover, multiple PNEs may exist in some types of games, but the price of anarchy only studies the deterioration in the worst PNE. Then Anshelevich et al. [3] introduced the term price of stability to refer to the ratio of the cost of best Nash equilibrium to the optimum network cost. In [3] they proved that a Shapley network design game with $n$ players always exists a PNE to which the price of stability with respect at most $H(n)=\sum^{n}_{i=1}\frac{1}{i}$ by using potential function introduced by Monderer and Shapley [9].
However, Anshelevich et al. [3] only proved few results when players are weighted, including proving the existence of PNE in all weighted Shapley network design game with two players. Chen and Roughgarden [7] presented a three-player weighted Shapley network design game that does not admit a PNE, which motivates subsequent study on $\alpha$-approximate pure Nash equlibrium ($\alpha$-APNE). In our model, $\alpha$-APNE is a profile that no players can unilaterally change strategy to reduce the cost by a factor of at least $\alpha$. Then Chen and Roughgarden [7] showed that the existence of $O(\log w_{max})$-APNE in every weighted Shapley network design game and proved the price of stability with respect to the $O(\log w_{max})$-APNE is $O(\log W)$, where $w_{max}$ and $W$ are the maximum weight and the sum of weights of all players respectively. In addition, they constructed a network without $o(\frac{\log w_{max}}{\log\log w_{max}})$-APNE.
On the other hand, computing a Nash equilibrium is essential when the game is given from a computational perspective. Apart from proposing Shapley network design game, Anshelevich et al. [3] showed that best response dynamics may take exponential time to converge to a PNE. Specifically, best response dynamics is an algorithm to find PNE ($\alpha$-APNE) by picking a player who has a strategy that can reduce personal cost (by more than $\alpha$ times) and updating the strategy as long as the profile is not a PNE ($\alpha$-APNE). Furthermore, Syrgkanis [11] showed that computing a PNE in the unweighted Shapley network design game is PLS-complete by the reduction from MAX CUT. Therefore, it is necessary to make further assumption on the game in order to ensure an algorithm that can find PNE or $\alpha$-APNE in a polynomial time. Giannakopoulos [12] assumed the constant cost $c_{e}$ of the edge as $\phi$-smooth independent random variables on $[0,1]$, which means its probability density function is upper bounded by $\phi$. And Giannakopoulos [12] proved that best response dynamics can compute the $\alpha$-APNE in polynomial time under this assumption.
In real situations, the cost of each edge in the network is not always constant, but changes with the number of players (or the sum of the weights of players) using the edge. However, to the best of our knowledge, no positive results exist regarding the weighted Shapley network design game where the cost of each edge is not a constant. Therefore, we extend the cost function of each edge to linear function to the total weights of the players that choose the edge. Under this assumption, we prove the existence of $\alpha$-APNE, put forward an algorithm to compute $\alpha$-APNE and analyze its time complexity.
@@ 1.1 Related Works
Epstein et al. [16] proved the existence of strong Nash equilibria of unweighted Shapley network design game, where strong Nash equilibria are profiles from which no group of players can decrease the cost of each of its members. Moreover, Epstein et al. [16] showed that the strong price of anarchy of a fair cost sharing game with $n$ players is at most $H(n)=\sum^{n}_{i=1}\frac{1}{i}$, namely $H(n)$ is the $n$-th harmonic number.
For undirected Shapley network design game with $n$ players, Anshelevich et al. [3] showed that the upper bound of the price of stability is $H(n)=\sum^{n}_{i=1}\frac{1}{i}$. Furthermore, Christodoulou et al. [1] present a lower bound of the price of stability of $42/23\approx 1.8261$, and was improved to $348/155\approx 2.245$ by Bilò et al. [15]. Fiat et al. [14] further restricted the network such that there is a common source vertex and a player in every vertex, and proved the price of stability of the game is $O(\log\log n)$. This further restricted game is called as broadcast network design game, and the price of stability of the game was improved to constant by Bilò et al. [10].
When the game does not apply the cost sharing mechanism, Fabrikant et al. [5] considered the game in which each player bears the whole cost of each chosen edge and named it the network congestion game. They proved that computing a Nash equilibrium is PLS-complete. Moreover, when all players have the same source vertex and sink vertex, namely the game is symmetric, Fabrikant et al. [5] showed that there is a polynomial algotithm for finding a PNE. Furthermore, Ackermann et al. [13] showed that the PLS-completeness holds in undirected network congestion game.
@@ 1.2 Our contributions
Our model is the first to extend the cost function to affine linear function in weighted Shapley network design game proposed by Anshelevich et al. [3], which is a constant $c_{e}$ for each edge $e$ in the network. In our model, the cost function of each edge in the network $c_{e}(w_{e})=a_{e}w_{e}+b_{e}$, where $a_{e},b_{e}>0$ are only related to the edge $e$ and $w_{e}$ is the total weight of the players that choose a path containing edge $e$. By constructing an $\alpha$-approximate potential function $\Phi(P)=\sum_{e\in E}c_{e}(w_{e})\log_{2}(1+w_{e})$, we prove the existence of $\alpha$-APNE, where $\alpha=2\log_{2}(1+W+w_{max})=O(\log_{2}(W))$, where $W$ and $w_{max}$ is the sum and maximum of the weight of all players respectively.
To compute the $\alpha$-APNE, we adopt $\epsilon$-best response dynamics which can obtain a $(1+\epsilon)\alpha$-APNE. Under the assumption that $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables, which means that the probability density function of $a_{e},b_{e}$ is upper bounded by a constant $\phi$, we prove the expected number of iterations is polynomial in $\frac{1}{\epsilon},\phi$, the number of players $n$ and the number of edges $m$ in the network. Specifically, we find out that the expected number of iterations can be bounded by $O(\frac{1}{\epsilon}W^{3}\phi n^{2}m\cdot ln(W\phi nm))$. Since in every step of iteration we can adopt the shortest-path search to verify the existence of $(1+\epsilon)\alpha$-improving deviation, then $\epsilon$-best response dynamics can compute the $(1+\epsilon)\alpha$-APNE in polynomial time.
Following the introduction, Section 2 defines the model and several concepts as preliminaries. Section 3 proves the existence of $\alpha$-APNE. Section 4 proposes an algorithm to compute the $\alpha$-APNE in polynomial time. Conclusions and possible directions in the future are presented in Section 5. An appendix gathers the mathematical techniques used in the main text.
@@ 2 Model and Preliminaries
We represent a weighted Shapley network design game $\Gamma$ by tuple $(\mathcal{G},\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ with components defined below.
  • P1: $\mathcal{G}=(V,E)$ is a directed graph, where $V$ is the vertex set of $\mathcal{G}$ and $E$ is the edge set of $\mathcal{G}$. Here we further define $m=|E|$, which is the number of edges in graph $\mathcal{G}$.
  • P2: $\mathcal{N}$ is a collection of $n$ different players $(n\geq 2)$. Moreover, each player $i\in\mathcal{N}$ is identified with a source-sink vertex pair $(s_{i},t_{i})$.
  • P3: $\mathcal{P}=(\mathcal{P}_{i})_{i\in\mathcal{N}}$, where $\mathcal{P}_{i}$ denote the set of simple $s_{i}-t_{i}$ paths. Furthermore, each player $i$ chooses a path $P_{i}\in\mathcal{P}_{i}$ from its source vertex to its sink vertex.
  • P4: $\mathcal{C}=(c_{e})_{e\in E}$, where $c_{e}(\cdot)$ is the non-negative cost function of the edge $e$.
@@
All players are selfish, and each of them chooses a path independently. Then it results in a profile $P=(P_{i})_{i\in\mathcal{N}}$, where $P_{i}$ denotes the simple $s_{i}-t_{i}$ path used by player $i\in\mathcal{N}$. For a given profile and an edge $e\in E$, we define $U_{e}(P):=\{i|i\in\mathcal{N},e\in P_{i}\}$ as the “user” of edge $e$. Moreover, we define $w_{e}=\sum_{i\in U_{e}(P)}w_{i}$ as the total weight of the players that choose a path containing edge $e$. Furthermore, for a given profile and a player $i$, the cost share $c^{i}_{e}$ of an edge $e\in E$ is $c_{e}(\cdot)\cdot w_{i}/w_{e}$. Then the cost to player $i$ in a profile is the sum of its cost shares: $c_{i}(P)=\sum_{e\in P_{i}}c^{i}_{e}$. In addition, the cost $c(P)$ of a profile $P$ is defined as $\sum_{i\in\mathcal{N}}c_{i}(P)$, which can be also written as $\sum_{e\in\cup_{i\in\mathcal{N}}P_{i}}c_{e}(\cdot)$.
@@
(pure Nash equilibrium) Given an instance $(\mathcal{G},\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game $\Gamma$. We call a profile $P=(P_{i})_{i\in\mathcal{N}}$ a pure Nash equilibrium(PNE) if, for each $i$, $P_{i}$ minimizes $c_{i}(P)$ over all paths in $\mathcal{P}_{i}$ while keeping $P_{j}$ fixed for $j\neq i$. That is:
@@
(exact potential function) Given an instance $(\mathcal{G},\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game $\Gamma$. We call a real-valued function $\Phi:\prod_{i\in\mathcal{N}}\mathcal{P}_{i}\rightarrow[0,\infty)$ an (exact) potential function if
@@
However, in some cases, the game may not admit a PNE. Then we have to consider a relaxed notion of $\alpha$-approximate pure Nash equilibrium instead.
@@
($\alpha$-approximate pure Nash equilibrium) Given an arbitrary instance $(\mathcal{G},\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game $\Gamma$ and an arbitrary constant $\alpha\geq 1$. We call a profile $P=(P_{i})_{i\in\mathcal{N}}$ an $\alpha$-approximate pure Nash equilibrium ($\alpha$-APNE) if
@@
In fact, Inequality (3) states that a unilateral change of strategy reduces the cost at a rate of at most $\frac{c_{i}(P_{i},P_{-i})-c_{i}(P^{\prime}_{i},P_{-i})}{c_{i}(P_{i},P_{-i})}\leq 1-\frac{1}{\alpha}$ in an $\alpha$-APNE. Moreover, we call a deviation $\alpha$-improving if the deviation decreases the cost incurred by the player by at least an $\alpha$ multiplicative factor. Therefore, $\alpha$-APNE is a profile from which no $\alpha$-improving deviations exist. We further point out that the smaller the constant $\alpha$ is, the closer the profile $P$ would approximate a PNE.
@@
($\alpha$-approximate potential function) Given an instance $(\mathcal{G},\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game $\Gamma$. For any player $i\in\mathcal{N}$, $P_{i},P^{\prime}_{i}\in\mathcal{P}_{i}$ satisfies $P_{i}\rightarrow P^{\prime}_{i}$ is an $\alpha$-improving deviation. We call a real-valued function $\Phi:\prod_{i\in\mathcal{N}}\mathcal{P}_{i}\rightarrow[0,\infty)$ an $\alpha$-approximate potential function if
@@ 3 Existence of $\alpha$-APNE
Anshelevich et al. [3] proved that the unweighted Shapley network design game admits a PNE, and pointed out that weighted Shapley network design game does not admit a PNE. Since our model is an extension of weighted Shapley network design game, we have to turn our attention to $\alpha$-APNE instead of PNE.
@@
Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game $\Gamma$. If the cost function is affine linear, namely $c_{e}(w_{e})=a_{e}w_{e}+b_{e}$, where $a_{e},b_{e}>0$ are only related to the edge $e$ and $w_{e}$ is the total weight of the players that choose a path containing edge $e$, then the game admits an $\alpha$-APNE, where $\alpha=O(\log_{2}(W))$.
Proof
For a weighted Shapley network design game $\Gamma$, we define a function $\Phi$ as follows:
$\Phi(P)=\sum_{e\in E}c_{e}(w_{e})\log_{2}(1+w_{e}).$
Now we consider an $\alpha$-improving deviation of player $i$ from the profile $P=(P_{i},P_{-i})$ to the profile $P^{\prime}=(P^{\prime}_{i},P_{-i})$. Without loss of generality, we can assume that $P_{i}$ and $P^{\prime}_{i}$ are disjoint. Otherwise, we can replace $P_{i}$ and $P^{\prime}_{i}$ with $P_{i}\backslash P^{\prime}_{i}$ and $P^{\prime}_{i}\backslash P_{i}$ in the following discussion. By the definition of $\alpha$-improving, we have
$\alpha\cdot\sum_{e\in P^{\prime}_{i}}{c_{e}(w_{e}+w_{i})}\cdot\frac{w_{i}}{w_{e}+w_{i}}\leq\sum_{e\in P_{i}}{c_{e}(w_{e})}\cdot\frac{w_{i}}{w_{e}}.$
Here, $w_{e}$ denotes the total weight on edge $e$ before the deviation and we let $\alpha=2\log_{2}(1+W+w_{max})=O(\log_{2}(W))$. Since $c_{e}(w_{e})=a_{e}w_{e}+b_{e}$, we have
$\alpha\cdot\sum_{e\in P^{\prime}_{i}}{[a_{e}(w_{e}+w_{i})+b_{e}]}\cdot\frac{w_{i}}{w_{e}+w_{i}}\leq\sum_{e\in P_{i}}{(a_{e}w_{e}+b_{e})}\cdot\frac{w_{i}}{w_{e}}.$ (5)
Then we can derive the following:
$\Delta\Phi$ $=\sum_{e\in P^{\prime}_{i}}{(a_{e}w_{e}+b_{e})}[\log_{2}(1+w_{e}+w_{i})-\log_{2}(1+w_{e})]$
$\,\quad-\sum_{e\in P_{i}}{(a_{e}w_{e}+b_{e})}[\log_{2}(1+w_{e})-\log_{2}(1+w_{e}-w_{i})]$
$\,\quad+\sum_{e\in P^{\prime}_{i}}a_{e}w_{i}\log_{2}(1+w_{e}+w_{i})-\sum_{e\in P_{i}}a_{e}w_{i}\log_{2}(1+w_{e}-w_{i})$
$<\sum_{e\in P^{\prime}_{i}}{(a_{e}w_{e}+b_{e})}\log_{2}[e(1+w_{i})]\frac{w_{i}}{w_{e}+w_{i}}$
$\,\quad-\sum_{e\in P_{i}}{(a_{e}w_{e}+b_{e})}\frac{w_{i}}{w_{e}}+\sum_{e\in P^{\prime}_{i}}a_{e}w_{i}\log_{2}(1+w_{e}+w_{i})$
$=\sum_{e\in P^{\prime}_{i}}a_{e}w_{i}\{\log_{2}(1+w_{e}+w_{i})+\log_{2}[e(1+w_{i})]\frac{w_{e}}{w_{e}+w_{i}}\}$
$\,\quad+\sum_{e\in P^{\prime}_{i}}b_{e}\log_{2}[e(1+w_{i})]\frac{w_{i}}{w_{e}+w_{i}}-\sum_{e\in P_{i}}{(a_{e}w_{e}+b_{e})}\frac{w_{i}}{w_{e}}$
$<\alpha\sum_{e\in P^{\prime}_{i}}{[a_{e}(w_{e}+w_{i})+b_{e}]}\cdot\frac{w_{i}}{w_{e}+w_{i}}-\sum_{e\in P_{i}}{(a_{e}w_{e}+b_{e})}\cdot\frac{w_{i}}{w_{e}}$
$=\alpha c_{i}(P^{\prime})-c_{i}(P).$
For the first inequality, we take the logarithm of $e(1+w_{i})\geq(1+\frac{w_{i}}{w_{e}+1})^{\frac{w_{e}+1}{w_{i}}}(1+\frac{w_{i}}{w_{e}+1})\geq(1+\frac{w_{i}}{w_{e}+1})^{\frac{w_{e}+w_{i}}{w_{i}}}$ and apply $\log_{2}(\frac{1+w_{e}}{1+w_{e}-w_{i}})>1\geq\frac{w_{i}}{w_{e}}$ when $2w_{i}\geq 1+w_{e}$ or take the logarithm of $(1+\frac{w_{i}}{1+w_{e}-w_{i}})^{\frac{w_{e}}{w_{i}}}\geq(1+\frac{w_{i}}{1+w_{e}-w_{i}})^{\frac{1+w_{e}-w_{i}}{w_{i}}}\geq 2$ when $2w_{i}<1+w_{e}$. Moreover, we use $\alpha\geq\max\{\log_{2}[e(1+w_{i})],\log_{2}(1+w_{e}+w_{i})+\log_{2}[e(1+w_{i})]\frac{w_{e}}{w_{e}+w_{i}}\}$ for the last inequality. Therefore, we have proved
$\Phi(P^{\prime})-\Phi(P)<\alpha c_{i}(P^{\prime})-c_{i}(P),$ (6)
which suggests that $\alpha$-improving deviations strictly decrease the function $\Phi$. Since the number of different profiles are finite, the theorem can be directly derived.
4 Compute the $(1+\epsilon)\alpha$-APNE
Initially we propose an algorithm to find the $(1+\epsilon)\alpha$-APNE of the game $\Gamma$ as followings, where $\epsilon$ is a positive constant close to 0.
Algorithm 1 $\epsilon$-approximate best response dynamics($\epsilon$-ABRD)
Input:
A weighted Shapley network design game $\Gamma=(\mathcal{G},\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$; a strategy profile $P\in\mathcal{P}$
Output:
A $(1+\epsilon)\alpha$-APNE of game $\Gamma$